# Group-robust Machine Unlearning

### v1

@@ C.3 Multi-group unlearning
This section further expands the experimental protocol of Sec. 4.4, highlighting the sampling composition and reporting all investigated metrics. For the 9 groups experiment we sampled the forget set data with either 20-29 y.o., 50-59 y.o., and 3-9 y.o. target attributes and afro-american, latino-hispanic, and white sensitive attributes. Instead, for the 25 groups experiment, we additionally sample from the more than 70 y.o. and 30-39 y.o. target attributes, and east asian and middle eastern sensitive attributes. The choice of groups is random and the unlearning ratio is fixed to 0.5.
@@
Full results for unlearning multiple groups in group-robust unlearning are reported in Tabs. 14 and 15. We notice overall the same trend as in Fig. 6. reweight contribution gets less important as the number of groups from which the forget set is sampled decreases, highlighted by unchanged UA, EO, and GA. The GA when using MIU, e.g., increases by nearly 5% in the 9 groups experiment (Tab. 14), while it remains unchanged in the 25 groups one (Tab. 15). Test accuracies are better preserved when unlearning 1 or 9 groups, while we notice a drop (about 1.5%) in the 25 groups experiment. We argue that this decline is caused by the larger forget set, which reduces the number of available samples in the retain set. Overall, MIU approximates Retrain + reweight better than baselines, consistently achieving the best Avg. Gap.
C.4 Extended ablation study
@@
Section 4.5 shows a comprehensive ablation of MIU’s components. However, Tab. 4 limits the analysis only to the UA, GA, and Avg. Gap to reduce space usage. Thus, Tab. 17 reports all metrics investigated for completeness. Although Tab. 4 metrics are limited, we chose a subset that shows great variance along different components and is more interesting to evaluate. For instance, EO variations are more nuanced compared to GA (e.g., it drops by 4.5 in FairFace [24], while GA grows by 11.1). The MIA shows subtle oscillations in CelebA [31] and FairFace [24] while it drops when adding the calibration term in Waterbirds [42], getting closer to Retrain + reweight value (59.5% of MIU vs. 53.6% of Retrain + reweight). Finally, TA and RA are relatively stable across components, with RA showing a negative trend (64.7% using all three components vs. 65.6% of Pretrain).
@@
Fig. 8, instead, reports the ablation study on parameter . We show results for and average each experiment over three different seeds. Overall, all three datasets benefit from the calibration term. However, while CelebA [31] and FairFace [24] achieve better results when (i.e., an Avg. Gap of 99.2 and 98.7), Waterbirds [42] minimizes the gap when (i.e., an Avg. Gap of 96.9). Nonetheless, results are generally stable, and when in doubt we suggest setting as it always shows an alignment improvement with Retrain + reweight.
C.5 Additional Fairness Metrics
@@
| CelebA [31] | ||||
| Pretrain | 44.1 | 21.3 | 20.9 | 65.5 |
| Retrain | 52.3 (8.2) | 33.9 (12.6) | 31.5 (10.6) | 55.6 (-9.9) |
| Retrain+rw | 43.7 (-0.4) | 21.0 (-0.3) | 20.8 (-0.1) | 66.7 (1.2) |
| MIU | 43.4 (-0.7) | 19.9 (-1.4) | 20.2 (-0.7) | 67.8 (2.3) |
| Waterbirds [42] | ||||
@@
| Dataset | Eq. 5 | Eq. 3 | Eq. 4 | RW | RA | UA | TA | MIA | EO | GA | Avg. Gap |
| CelebA [31] | 85.10.0 | 53.60.7 | 82.60.1 | 0.20.0 | 28.60.1 | 54.01.0 | 94.10.6 | ||||
| 84.80.0 | 55.30.8 | 82.60.2 | 0.30.1 | 27.40.4 | 55.91.0 | 94.90.7 | |||||
| 63.17.1 | 41.734.2 | 62.17.0 | 0.00.0 | 10.45.4 | 42.334.1 | 78.98.8 | |||||
| 84.20.1 | 68.80.3 | 82.50.1 | 0.10.0 | 20.20.6 | 69.01.2 | 99.20.6 | |||||
| Waterbirds [42] | 100.00.0 | 47.67.3 | 85.00.6 | 73.83.4 | 32.30.8 | 51.10.6 | 92.54.6 | ||||
| FairFace [24] | 65.20.1 | 63.11.6 | 56.90.3 | 0.30.0 | 10.30.8 | 59.21.9 | 96.10.8 | ||||
| 66.70.2 | 74.71.2 | 57.20.7 | 0.30.0 | 6.02.0 | 66.14.4 | 98.10.4 | |||||
| 59.13.1 | 87.16.8 | 54.51.8 | 0.00.0 | 3.10.5 | 81.16.2 | 93.03.1 | |||||
| 64.70.3 | 71.62.8 | 57.10.3 | 0.30.2 | 5.80.4 | 70.31.6 | 98.70.8 |
To further study the method’s fairness after unlearning. We investigate three additional fairness metrics alongside the Equalized Odds (EO) [16], namely, Demographic Parity [29], Equal Opportunity [16], and Worst Group Accuracy (WG) [42, 30]. To satisfy the Demographic Parity, the model probability of outputting a positive prediction must be independent of the sensitive attribute. Therefore, we measure it as: . Similarly, to satisfy Equal Opportunity (EO), the model true positive rate must be independent of the sensitive attribute: . Instead, the Worst Group Accuracy measures the average accuracy of the worst group of the test set. Thus, the Worst Group Accuracy is computed as: , where is the subset of images of group of the test set. The lower the first two metrics (DP and EP) the better the model fairness, while the higher the WG the higher the model robustness.
Table 16 reports our evaluation using the additional metrics on CelebA [31], Waterbirds [42], and FairFace [24]. While plain model retraining generally shows performance degradation on all fairness metrics (e.g., +8.2 DP, +12.6 EP, +10.6 EO, and -9.9 WG, in CelebA [31]), using reweight recovers the original model fairness and overall robustness (i.e., -0.4 DP, -0.3 EP, -0.1 EO, and +1.2 WG, in CelebA [31]). Similarly, the proposed MIU preserves Pretrain robustness by approximating Retrain + reweight (i.e., -0.7 DP, -1.4 EP, -0.7 EO, and +2.3 WG, in CelebA [31]).
Appendix D Limitations and negative societal impacts
Limitations. One limitation of our work is assuming knowledge of the data point group. Group annotations are difficult to obtain. Therefore, a natural follow-up for this work is a group-agnostic methodology that preserves model robustness like reweight. Furthermore, our evaluation is limited to the image classification task. Using the proposed techniques and baselines in different scenarios might be non-trivial and there is no guarantee that unlearning effectiveness and accuracy will be preserved. Nevertheless, our work is the first step toward quantifying and mitigating the accuracy degradation caused by non-uniformly sampled forget sets.
Negative societal impacts. Machine unlearning is designed to remove user data from a trained model. While its primary goal is privacy preservation, it can also be misused for malicious purposes. Targeted unlearning of specific groups may lead to biased models with harmful consequences. However, the proposed group-robust machine unlearning seeks to minimize performance degradation for dominant groups in the forget set. Thus, MIU and reweight counteract biases introduced by unlearning, mitigating the negative societal impacts of its misuse.

### v2

@@ B.3 Multi-group unlearning
This section further expands the experimental protocol of Sec.˜4.4, highlighting the sampling composition and reporting all investigated metrics. For the 9 groups experiment, we sampled the forget set data with either 20-29 y.o., 50-59 y.o., and 3-9 y.o. target attributes, and Afro-American, Latino-Hispanic, and Caucasian sensitive attributes. Instead, for the 25 groups experiment, we additionally sample from the more than 70 y.o. and 30-39 y.o. target attributes, and East Asian and Middle Eastern sensitive attributes. The choice of groups is random, and the unlearning ratio is fixed to 0.5.
@@
Full results for unlearning multiple groups in group-robust unlearning are reported in Tabs.˜14 and 15. We notice overall the same trend as in Fig.˜6. reweight contribution gets less important as the number of groups from which the forget set is sampled decreases, highlighted by unchanged UA, EO, and GA. The GA, when using MIU, e.g., increases by nearly 5% in the 9 groups experiment (Tab.˜14), while it remains unchanged in the 25 groups one (Tab.˜15). Test accuracies are better preserved when unlearning 1 or 9 groups, while we notice a drop (about 1.5%) in the 25 groups experiment. We argue that this decline is caused by the larger forget set, which reduces the number of available samples in the remaining set. Overall, MIU approximates Retrain + reweight better than baselines, consistently achieving the best avg. gap.
B.4 Unlearning on MultiNLI
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
| Pretrain | 94.72.5 | 81.88.4 | 82.30.1 | 31.40.9 | 14.41.8 | 59.32.7 | - | |
| Retrain | 95.91.1 | 53.03.7 | 82.30.1 | 47.63.9 | 19.30.6 | 50.83.1 | - | |
| Retrain | 95.31.1 | 59.80.7 | 81.70.0 | 43.32.3 | 16.20.3 | 58.82.0 | - | |
| L1-sparse (Jia et al., 2023) | 84.82.0 | 46.64.8 | 79.00.7 | 34.12.3 | 23.31.4 | 42.02.1 | 90.11.8 | |
| SalUn (Fan et al., 2024b) | 76.430.3 | 36.525.9 | 63.821.4 | 36.125.6 | 20.20.8 | 35.124.9 | 80.819.0 | |
| SCRUB (Kurmanji et al., 2024) | 96.41.9 | 56.84.6 | 82.40.2 | 46.75.5 | 22.61.7 | 47.63.3 | 95.31.3 | |
| MIU | 99.20.1 | 60.31.6 | 79.20.2 | 54.92.8 | 18.80.7 | 55.62.0 | 95.80.3 | |
| L1-sparse (Jia et al., 2023) | 84.92.0 | 55.43.8 | 79.00.6 | 31.71.4 | 19.31.1 | 49.21.6 | 93.01.2 | |
| SalUn (Fan et al., 2024b) | 96.00.1 | 54.33.6 | 78.30.2 | 48.12.0 | 18.51.9 | 51.92.6 | 95.92.3 | |
| SCRUB (Kurmanji et al., 2024) | 96.51.9 | 82.17.8 | 82.40.2 | 32.28.9 | 14.03.8 | 62.46.3 | 92.52.8 | |
| MIU | 99.90.0 | 67.02.6 | 81.50.2 | 51.12.6 | 18.00.4 | 56.61.2 | 96.00.8 |
Table˜16 shows the result for group-robust unlearning on MultiNLI (Williams et al., 2018) using an unlearning ratio of 0.5. reweight allows a successful recovery of the original group accuracy (-0.5), while preserving the test accuracy (-0.6). Compared to Pretrain, the EO are worse (14.4 vs. 16.2) but better compared to plain Retrain (16.2 vs.19.3), thus, reweight partly recovered also the equalized odds. MIU achieves the best performance preservation overall by scoring the highest GA alignment, both when using (96.0) and not using (95.8) reweight. Particularly, MIU is less influenced by reweight, and most of the alignment comes from its unique loss formulation. SCRUB (Kurmanji et al., 2024) achieves the second-best result without reweight (95.3), but struggles when used (92.5), as forget accuracy unexpectedly grows to high values (82.1). Instead, SalUn (Fan et al., 2024b) scores the best avg. gap after MIU (95.9), but shows high variance and low results without reweight. Overall, MIU is the most consistent among the two configurations and generally shows the lowest variance across metrics.
| dataset | Eq.˜7 | Eq.˜3 | Eq.˜6 | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
| CelebA | 96.40.1 | 35.50.9 | 95.90.0 | 1.10.1 | 26.70.9 | 35.42.3 | 97.50.5 | ||||
| 96.40.1 | 36.30.9 | 95.90.1 | 1.00.3 | 26.10.8 | 36.32.0 | 98.00.4 | |||||
| 95.80.5 | 27.85.7 | 95.30.4 | 0.90.2 | 25.50.8 | 28.56.9 | 95.22.5 | |||||
| 96.30.2 | 43.20.6 | 96.00.0 | 1.20.1 | 24.00.5 | 41.31.1 | 99.00.2 | |||||
| Waterbirds | 100.00.0 | 47.67.3 | 85.00.6 | 73.83.4 | 32.30.8 | 51.10.6 | 92.54.6 | ||||
| FairFace | 65.20.1 | 63.11.6 | 56.90.3 | 0.30.0 | 10.30.8 | 59.21.9 | 96.10.8 | ||||
| 66.70.2 | 74.71.2 | 57.20.7 | 0.30.0 | 6.02.0 | 66.14.4 | 98.10.4 | |||||
| 59.13.1 | 87.16.8 | 54.51.8 | 0.00.0 | 3.10.5 | 81.16.2 | 93.03.1 | |||||
| 64.70.3 | 71.62.8 | 57.10.3 | 0.30.2 | 5.80.4 | 70.31.6 | 98.70.8 |
B.5 Extended ablation study
@@
Section˜4.5 shows a comprehensive ablation of MIU’s components. However, Tab.˜4 limits the analysis only to the UA, GA, and avg. gap to reduce space usage. Thus, Tab.˜17 reports all metrics investigated for completeness. Although Tab.˜4 metrics are limited, we chose a subset that shows great variance along different components and is more interesting to evaluate. For instance, EO variations are more nuanced compared to GA (e.g., it drops by 4.5 in FairFace (Karkkainen & Joo, 2021), while GA grows by 11.1).
The MIA shows subtle oscillations in CelebA (Liu et al., 2015) and FairFace (Karkkainen & Joo, 2021) while it drops when adding the calibration term in Waterbirds (Sagawa et al., 2020), getting closer to Retrain + reweight value (59.5% of MIU vs. 53.6% of Retrain + reweight). Finally, TA and RA are relatively stable across components, with RA showing a negative trend (64.7% using all three components vs. 65.6% of Pretrain).
@@
Fig.˜8, instead, reports the ablation study on parameter . We show results for and average each experiment over three different seeds. Overall, all three datasets benefit from the calibration term. However, while CelebA (Liu et al., 2015) and FairFace (Karkkainen & Joo, 2021) achieve better results when (i.e., an avg. gap of 99.2 and 98.7), Waterbirds (Sagawa et al., 2020) minimizes the gap when (i.e., an avg. gap of 96.9). Nonetheless, results are generally stable, and when in doubt, we suggest setting as it always shows an alignment improvement with Retrain + reweight.
@@
| CelebA (Liu et al., 2015) | ||||
| Pretrain | 18.7 | 45.3 | 24.2 | 40.6 |
| Retrain | 18.9 (0.1) | 51.0 (5.7) | 27.0 (2.9) | 34.4 (-6.1) |
| Retrain+rw | 18.8 (0.1) | 44.6 (-0.7) | 23.9 (-0.3) | 41.3 (0.7) |
| MIU | 18.8 (0.1) | 45.0 (-0.3) | 24.0 (-0.1) | 41.3 (0.7) |
| Waterbirds (Sagawa et al., 2020) | ||||
@@
B.6 Additional Fairness Metrics
To further study the method’s fairness after unlearning. We investigate three additional fairness metrics alongside the Equalized Odds (EO) (Hardt et al., 2016), namely, Demographic Parity (Kusner et al., 2017), Equal Opportunity (Hardt et al., 2016), and Worst Group Accuracy (WG) (Sagawa et al., 2020; Liu et al., 2021). To satisfy the Demographic Parity, the model’s probability of outputting a positive prediction must be independent of the sensitive attribute. Therefore, we measure it as: . Similarly, to satisfy Equal Opportunity (EP), the model true positive rate must be independent of the sensitive attribute: . Instead, the Worst Group Accuracy measures the average accuracy of the worst group of the test set. Thus, the Worst Group Accuracy is computed as: , where is the subset of images of group of the test set. The lower the first two metrics (DP and EP), the better the model fairness, while the higher the WG, the higher the model robustness.
Table˜18 reports our evaluation using the additional metrics on CelebA (Liu et al., 2015), Waterbirds (Sagawa et al., 2020), and FairFace (Karkkainen & Joo, 2021). While plain model retraining generally shows performance degradation on almost all fairness metrics (e.g., +5.7 EP, +2.9 EO, and -6.1 WG, in CelebA (Liu et al., 2015)), using reweight recovers the original model fairness and overall robustness (i.e., -0.7 EP, -0.3 EO, and +0.7 WG, in CelebA (Liu et al., 2015)). Similarly, the proposed MIU preserves Pretrain robustness by approximating Retrain + reweight (i.e., -0.3 EP, -0.1 EO, and +0.7 WG, in CelebA (Liu et al., 2015)).